OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80mph and train B is traveling at 100mph. Train A passes a station at 9:20pm if train B passes the same station at 9:35 PM, at what time will train B catch up to train A?

OpenStudy (radar):

use the d=rt formula distance = rate x time. When both trains have traveled the same distance from the station they will be at the same point and caught up with one another. d for train A = (80 mph)(t) d for train B =100 mph(t-15 min (this train passed station 15 minutes later than A) since the distance is the same we can write: 80t=100t -1500 -20t=-1500 t=75 minutes time 9:20 + 75 min=10:35 to verify: train a travels 1.25 hr at 80 mph for a distance of 100 miles train b travels 1.0 hr at 100 mph for a distance of 100 miles,yup further clarification: 75 minutes = 1.25 hr train b time was 75-15 minutes = 60 min or 1 hr.

OpenStudy (radar):

10:35 PM of course.